OpenStudy (anonymous):

Write the equivalent rates. 3/8=__/2 = __/12 = __/20

OpenStudy (anonymous):

3/8=1.5/4=.75/2

OpenStudy (anonymous):

1.5/4=4.5/12 (you multiply both top and bottom by 3)

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

1.5/4=7.5/20 (multiply both top and bottom by 5)

OpenStudy (anonymous):

idk what you keep talking about

OpenStudy (anonymous):

3/8=.75/2=4.5/12=7.5/20 is the answer

OpenStudy (anonymous):

ok thx LOL